Overtime function:
 You have a job where you make $10 per hour.
 After you work 40 hours you get paid time and one half.
 You may only work a maximum of 60 hours.
Write a piecewise function that represents your pay for up to 60 hours.

Answer by Theo(13342)

y = 10 * x for x <= 40
y = 400 + 15 * (x - 40) for 40 < x <= 60

if you worked less than or equal to 40 hours, your pay is 10 * the number of hours you worked.

if you worked more than 40 hours and less than or equal to 60 hours, your pay is 400 + 15 * (the number of hours worked minus 40).

you worked 35 hours.
your pay is 10 * 35 = 350 dollars.

you worked 47 hours.
your pay is 400 + (47 - 40) = 7 * 15 = 400 + 105 = 505 dollars.
note that the 400 dollars covers the first 40 hours at 10 dollars an hour.

you worked 60 hours.
your pay is 400 + (60 - 40) * 15 = + 20 * 15 = 300 = 700.
note that the 400 dollars covers the first 40 hours at 10 dollars an hour.
then you get 15 dollars for the next 20 hours.